# SQL-R1: Training Natural Language to SQL Reasoning Model By Reinforcement Learning

### v1

@@ Abstract
Natural Language to SQL (NL2SQL) enables intuitive interactions with databases by transforming natural language queries into structured SQL statements. Despite recent advancements in enhancing human-computer interaction within database applications, significant challenges persist, particularly regarding the inference performance in complex scenarios involving multi-table joins and nested queries. Current methodologies primarily utilize supervised fine-tuning (SFT) to train the NL2SQL model, which may limit adaptability and interpretability in new environments  (e.g., finance and healthcare). In order to enhance the reasoning performance of the NL2SQL model in the above complex situations, we introduce SQL-R1, a novel NL2SQL reasoning model trained by the reinforcement learning (RL) algorithms. We design a specialized RL-based reward function tailored for NL2SQL tasks and discussed the impact of cold start on the effectiveness of intensive training. In addition, we achieve competitive accuracy using only a tiny amount of synthetic NL2SQL data for augmented training and further explore data engineering for RL. In existing experiments, SQL-R1 achieves execution accuracy of 88.6% and 66.6% on the benchmark Spider and BIRD, respectively, only using the 7B base model.
@@ 1 Introduction
Despite these advancements, improving the NL2SQL system’s inference performance in complex database scenarios remains a considerable challenge. As shown in Figure 1, schema complexity may lead to generation errors in processing multi-table joins and nested queries, and it is difficult for individually trained models to think and process complex semantics. Currently, a significant portion of NL2SQL research is devoted to training open-source large language models (LLMs) by supervised fine-tuning (SFT) (dtssql, ; codes, ; catsql, ) to achieve accuracy at a smaller model scale compared to approaches using closed-source LLMs (e.g., GPT-4, GPT-4o) (dailsql, ; dinsql, ; c3sql, ). However, SFT relies on the database schema’s structure and the training data scale. This may lead to the existing model’s instability in domain adaptation and generalization in new database environments. Additionally, the lack of interpretability of NL2SQL reasoning logic limits the application of the model in high-risk fields, such as finance and healthcare.
@@
Recently, reinforcement learning (RL) has shown great potential in training the reasoning ability of LLMs in recent research. Compared with supervised fine-tuning, reinforcement learning can dynamically adjust the decision-making strategy of the LLMs through interaction with the environment, thereby achieving superior performance in complex reasoning tasks (deepseekr1, ). RL-based methods have proven effective in enhancing model reasoning and generalization capabilities in financial reasoning (finr1, ), search engines (searchr1, ) and mathematical reasoning (logicrl, ; openr1, ).
@@
  • Explicit NL2SQL Reasoning Model: We propose SQL-R1, a NL2SQL reasoning model trained on a few NL2SQL data (e.g., 5K) currently, which can achieve 88.6% and 66.6% accuracy on the leading benchmark Spider-Test and BIRD respectively, and can output detailed explicit reasoning process.
  • Training Strategy for NL2SQL Reasoning Model: We explore the impact of cold-start training on the performance of SQL-R1 and construct a training strategy based on SFT and RL, which can effectively improve the NL2SQL inference performance of the model.
@@ 2.3.3 Reward Function Design
$S_{f}=\begin{cases}1,&\text{if format is correct}\\ -1,&\text{if format is incorrect}\end{cases}$
Execution Reward
@@
$S_{e}=\begin{cases}2,&\text{if SQL candidate is executable}\\ 0,&\text{if format is incorrect}\\ -2,&\text{if SQL candidate is not executable}\end{cases}$
Result Reward
@@
$S_{r}=\begin{cases}3,&\text{if query result is correct}\\ 0,&\text{if format is incorrect or SQL candidate is not executable}\\ -3,&\text{if query result is incorrect}\end{cases}$
Length Reward
@@
$S_{l}=\begin{cases}0.5\times S_{tl}+S_{al},&\text{if query result is correct and $len_{response}<=$ {MAX LENGTH}}\\ 0.5+S_{al},&\text{if query result is correct and $len_{response}>$ {MAX LENGTH}}\\ 0,&\text{other cases}\\ \end{cases}$
@@ 3.1 Setup
Implementation Settings
Currently, SQL-R1 is mainly built on Qwen2.5-Coder-7B-Instruct. For the SFT training, we set the learning rate as 5e-5; batch size as 1. For the RL training, we set the learning rate as 3e-7, rollout of actor model as 8; max response length as 2048. For inference, we set the count of SQL candidates as 8 and the temperature as 0.8.
@@ 3.2 Main Results
Table 1: Execution accuracy of different NL2SQL methods on Spider and BIRD benchmark.
@@
OmniSQL (omnisql, ) Qwen2.5-Coder-7B Self-Consistency 81.6 88.9 66.1
@@
Performance on Main Benchmarks
The results presented in Table 1 highlight the remarkable performance of SQL-R1. It managed to attain the best execution accuracy of 87.6% in the Spider development dataset, 88.7% in the test dataset, and 66.6% in the BIRD development dataset. Compared to the NL2SQL solution that employs the Qwen2.5-Coder-7B, our proposed model demonstrates comparable performance levels. Notably, SQL-R1 exhibits superior performance relative to numerous NL2SQL solutions that rely on closed-source models (e.g. GPT-4 and GPT-4o).
Insights of Performance and Model Scale Trade-offs
As illustrated in Figure 2, we investigated the relationship between performance and model size utilizing on BIRD development dataset. The comparative analysis encompasses various model types, including NL2SQL models, reasoning LLMs and general LLMs. For GPT-4, GPT-4o, GPT-4o-mini and Gemni-1.5-Pro, we refer to the parametric description of (abacha2024medec, ; omnisql, ; alphasql, ). The findings demonstrate that when utilizing the 7B model as a foundational model, SQL-R1 attains accuracy levels that surpass those of larger-scale models, particularly. This underscores the efficacy of the proposed model in optimizing the performance of NL2SQL while ensuring that cost efficiency is maintained.
Case Study
@@ 3.3 Analysis of SFT Cold Start
In this experiment, we adopt three different baselines for comparison, namely the initial model that has not been pretrained, the model that uses the original format for SFT (OmniSQL-7B), and the model that uses the reinforced instruction format for SFT. As demonstrated in Table 2, cold-start training significantly enhances the performance of RL-based NL2SQL reasoning models. We observed improved instruction-following capabilities and SQL generation quality during the RL phase by implementing two forms of cold-start strategies. For Q2, these result confirms the necessity of cold-start training to optimize the model’s exploration and reasoning performance in complex database scenarios. At present, we are still conducting further experiments, and relevant results will continue to be updated in subsequent versions.
For Q3, we currently employ a small-scale subset of the SynSQL-2.5M dataset for the purpose of reinforcement training. This methodology demonstrates substantial advancements within the models, thereby showcasing the viability of utilizing synthetic data engineering to augment model inference capabilities in the NL2SQL domain. Subsequent publications will outline further strategies pertaining to synthetic data engineering.
Table 2: Execution accuracy of models with different cold start strategy.
@@
SQL-R1 + Qwen2.5-Coder-7B
@@
SQL-R1 + OmniSQL-7B (omnisql, ) SynSQL-2.5M
@@ 3.4 Ablation Study of Reward Components
We conducted an ablation experiment on the BIRD development dataset to verify the proposed reinforcement learning reward mechanism’s effectiveness. The experiment sequentially removed individual reward components from the comprehensive reward function while maintaining the parameter settings established in Section 3.1. As presented in Table 3, the result demonstrates that omitting any reward component from the original reward function adversely impacts inference performance. This underscores the critical importance of execution feedback and result reward in the model training process.
@@ 4 Limitations
Supported Database Dialect
The current research only supports training and evaluation on the SQLite dialect, and no further evaluation has been done on more database dialects (e.g., Snowflake, DuckDB).
@@ Appendix A Detailed Case Study
Appendix B Prompt Templates for Training

### v2

@@ Abstract
Natural Language to SQL (NL2SQL) enables intuitive interactions with databases by transforming natural language queries into structured SQL statements. Despite recent advancements in enhancing human-computer interaction within database applications, significant challenges persist, particularly regarding the inference performance in complex scenarios involving multi-table joins and nested queries. Current methodologies primarily utilize supervised fine-tuning (SFT) to train the NL2SQL model, which may limit adaptability and interpretability in new environments  (e.g., finance and healthcare). In order to enhance the reasoning performance of the NL2SQL model in the above complex situations, we introduce SQL-R1, a novel NL2SQL reasoning model trained by the reinforcement learning (RL) algorithms. We design a specialized RL-based reward function tailored for NL2SQL tasks and discussed the impact of cold start on the effectiveness of intensive training. In addition, we achieve competitive accuracy using only a tiny amount of synthetic NL2SQL data for augmented training and further explore data engineering for RL. In existing experiments, SQL-R1 achieves execution accuracy of 88.6% and 67.1% on the benchmark Spider and BIRD, respectively.
@@ 1 Introduction
Despite these advancements, improving the NL2SQL system’s inference performance in complex database scenarios remains a considerable challenge. As shown in Figure 1, schema complexity may lead to generation errors in processing multi-table joins and nested queries, and it is difficult for individually trained models to think and process complex semantics. Currently, a significant portion of NL2SQL research is devoted to training open-source large language models (LLMs) by supervised fine-tuning (SFT) (dtssql, ; codes, ; catsql, ) to achieve accuracy at a smaller model scale compared to approaches using closed-source LLMs (e.g., GPT-4, GPT-4o) (dailsql, ; dinsql, ; c3sql, ). However, SFT relies on the database schema structure and the training data scale. This may lead to the existing model’s instability in domain adaptation and generalization in new database environments. Additionally, the lack of interpretability of NL2SQL reasoning logic limits the application of the model in high-risk fields, such as finance and healthcare.
@@
Recently, reinforcement learning (RL) has shown great potential in training the reasoning ability of LLMs in recent research. Compared with supervised fine-tuning, reinforcement learning can dynamically adjust the decision-making strategy of the LLMs through interaction with the environment, thereby achieving superior performance in complex reasoning tasks (deepseekr1, ). RL-based methods have proven effective in enhancing model reasoning and generalization capabilities in financial reasoning (finr1, ), search engines (searchr1, ), and mathematical reasoning (logicrl, ; openr1, ).
@@
  • Explicit NL2SQL Reasoning Model: We propose SQL-R1, a NL2SQL reasoning model trained on a few NL2SQL data (e.g., 5K) currently, which can achieve 88.6% and 66.6% accuracy on the leading benchmark Spider-Test and BIRD, respectively, and can output a detailed explicit reasoning process.
  • Training Strategy for NL2SQL Reasoning Model: We extensively explored the impact of cold-start training on SQL-R1, developing a training strategy that integrates SFT and RL. Our findings highlight the strategy of using synthetic data to enhance model performance and robustness, offering key insights into optimizing NL2SQL reasoning model training.
@@ 2.3.3 Reward Function Design
$S_{f}=\begin{cases}1,&\text{if format is correct}\\ -1,&\text{if format is incorrect}\end{cases}$ (2)
Execution Reward.
@@
$S_{e}=\begin{cases}2,&\text{if SQL candidate is executable}\\ 0,&\text{if format is incorrect}\\ -2,&\text{if SQL candidate is not executable}\end{cases}$ (3)
Result Reward.
@@
$S_{r}=\begin{cases}3,&\text{if query result is correct}\\ 0,&\text{if format is incorrect or SQL candidate is not executable}\\ -3,&\text{if query result is incorrect}\end{cases}$ (4)
Length Reward.
@@
$S_{l}=\begin{cases}0.5\times S_{tl}+S_{al},&\text{if query result is correct and $len_{response}<=$ {MAX LENGTH}}\\ 0.5+S_{al},&\text{if query result is correct and $len_{response}>$ {MAX LENGTH}}\\ 0,&\text{other cases}\\ \end{cases}$ (5)
@@ 3.1 Setup
Implementation Settings.
Currently, SQL-R1 is mainly built on Qwen2.5-Coder series models (qwen2.5, ). For the SFT training, we set the learning rate as 5e-5; batch size as 1. For the RL training, we set the learning rate as 3e-7, rollout of actor model as 8; max response length as 2048. For inference, we set the count of SQL candidates as 8 and the temperature as 0.8.
@@ 3.2 Main Results
Table 1: Execution accuracy (%) of different NL2SQL methods on Spider and BIRD benchmark.
@@
OmniSQL (omnisql, ) Qwen2.5-Coder-7B Self-Consistency 85.5 88.9 66.1
DeepRetrieval (deepretrieval, ) Qwen2.5-Coder-7B - - 76.1 56.0
Reasoning-SQL (reasoningsql, ) Qwen2.5-Coder-14B Self-Consistency 81.4 - 65.3
@@
CHASE-SQL (chasesql, ) Gemimi-1.5-Pro - - 87.6 73.0
SQL-R1 (Ours) Qwen2.5-Coder-3B Self-Consistency 78.1 78.9 54.6
@@
SQL-R1 (Ours) Qwen2.5-Coder-14B Self-Consistency 86.7 88.1 67.1
Performance on Main Benchmarks.
The results presented in Table 1 underscore the performance of SQL-R1 across various foundational models. Specifically, when trained on the Qwen2.5-Coder-3B, SQL-R1 achieved an execution accuracy of 78.1% on the Spider development dataset, 78.9% on the Spider test dataset, and 54.6% on the BIRD development dataset, demonstrating robust performance even with a smaller base model. Moreover, SQL-R1 exhibits a marked enhancement in accuracy when utilized with larger models; for instance, the Qwen2.5-Coder-7B model yielded accuracies of 87.6%, 88.7%, and 63.1% across the same Spider development, Spider test, and BIRD development datasets, respectively. The deployment of the even larger Qwen2.5-Coder-14B model further affirmed SQL-R1’s capabilities, achieving high accuracy rates of 86.7%, 88.1%, and 67.1% on these datasets. In comparison to other solutions leveraging closed-source models, such as GPT-4 and GPT-4o, SQL-R1 demonstrates competitive and often superior performance. Notably, In comparison to contemporaneous exploration of reasoning-based NL2SQL (e.g., DeepRetrieval (deepretrieval, ) and Reasoning-SQL (reasoningsql, )), SQL-R1 surpasses them on the Spider-Dev and BIRD-Dev datasets. This performance highlights the efficacy of our proposed approach in utilizing diverse base models to attain state-of-the-art results on complex NL2SQL reasoning tasks. Additional experimental results on other NL2SQL benchmarks can be found in Section B.1 and Table 4.
Insights of Performance and Model Scale Trade-offs.
As illustrated in Figure 2, we investigated the relationship between performance and model size utilizing on BIRD development dataset. The comparative analysis encompasses various model types, including NL2SQL models, reasoning LLMs and general LLMs. For GPT-4, GPT-4o, GPT-4o-mini and Gemni-1.5-Pro, we refer to the parametric description of (abacha2024medec, ; omnisql, ; alphasql, ). The findings demonstrate that when utilizing the 7B model as a foundational model, SQL-R1 attains accuracy levels that surpass those of larger-scale models, particularly. This underscores the efficacy of the proposed model in optimizing the performance of NL2SQL while ensuring that cost efficiency is maintained. Additionally, Section B.2 provides some detailed results of different base models.
Case Study.
@@ 3.3 Analysis of SFT Cold Start
Table 2: Execution accuracy (%) of models with different cold start strategy. The Reasoning Instruction column is applied represents the instruction applied SFT process.
@@
OmniSQL-14B (omnisql, ) SynSQL-2.5M
86.2 88.3 65.9
SQL-R1 + Qwen2.5-Coder-7B -
@@
SQL-R1 + Qwen2.5-Coder-14B -
86.7 88.1 67.1
SQL-R1 + OmniSQL-7B SynSQL-2.5M
@@
SQL-R1 + OmniSQL-14B SynSQL-2.5M
86.4 87.6 66.6
In this experiment, we conducted a comprehensive evaluation on different baselines: the base model without any post-training, the model utilizing the conventional format for SFT (e.g., OmniSQL-7B & OmniSQL-14B), and the model employing a reasoning instruction format for SFT.
Based on the findings demonstrated in Table 2, several implications can be discerned: Firstly, the quantity of synthesized data significantly influences the performance and generalization capabilities of SFT, as well as the effectiveness of subsequent RL training. For instance, the SQL-R1 7B model, which was initiated using SynSQL-2.5M, demonstrates superior performance compared to the model that trained with SynSQL-200K. Moreover, the provenance of the training data also plays a critical role in determining the performance of SFT. Notably, models that rely on synthetic data for SFT exhibited only marginal gains when evaluated on the BIRD-Dev dataset. Specifically, the SFT methodology utilizing synthetic data did not yield an improvement in accuracy for the 7B model, even when the inference format adhered to the established instructions and utilized a training set of 200K instances. Additionally, the SFT training results for the 14B models in both Reasoning-SQL and OmniSQL underscore the importance of using a training set derived from a consistent source for SFT.
Thus, in response to Q2, these findings imply that SFT cold-start training is not universally essential for RL-based NL2SQL models. Its effectiveness is contingent upon the origin and volume of the training data. Future investigations should focus on delineating the optimal conditions that enable cold-start training to confer significant benefits, particularly with respect to data source and volume.
In addressing Q3, experimental results corroborate the assertion that synthetic data engineering plays a pivotal role in augmenting the inference capabilities of NL2SQL models, especially when utilized in conjunction with RL. The synergistic effect of synthetic data and RL training reveals marked enhancements in model performance. For instance, SQL-R1, when trained on SynSQL-Complex-5K, surpasses the performance of both the 7B and 14B baseline models. This finding elucidates the substantial potential inherent in synthetic data engineering to enhance model capabilities. Ongoing research endeavors will persist in exploring and refining strategies for the integration of synthetic data into the training regimen, with the aim of further advancing model performance and generalization.
@@ 3.4 Ablation Study of Reward Components
We conducted an ablation experiment on the BIRD development dataset to verify the proposed reinforcement learning reward mechanism’s effectiveness. The experiment sequentially removed individual reward components from the comprehensive reward function while maintaining the parameter settings established in Section 3.1. As presented in Table 3, the result demonstrates that omitting any reward component from the original reward function adversely impacts inference performance. This underscores the critical importance of execution feedback and result reward in the model training process. Section B.3 and B.4 provide more details of the ablation study.
@@
4 Related Works
NL2SQL Methods.
Recent advancements in large language models have significantly propelled the development of NL2SQL translation techniques (DBLP:journals/corr/abs-2408-05109, ; katsogiannis2023survey, ; hong2024next, ). Current research focuses on optimizing various components of the NL2SQL workflow, including pre-processing modules like schema linking (resdsql, ), translation strategies (scprompt, ; zeronl2sql, ; natsql, ; wolfson2020break, ; eyal2023semantic, ; gao2022towards, ; petsql, ), and post-processing modules like SQL correction (liu2025nl2sql, ; dinsql, ; macsql, ) and self-consistency for SQL selection (dailsql, ; c3sql, ; omnisql, ). However, existing NL2SQL models primarily rely on supervised fine-tuning, which may exhibit instability in domain adaptation and new scenario fitting, as they heavily depend on the knowledge of database schema and structure from the training data (dtssql, ; catsql, ; ratsql, ). Additionally, the lack of interpretability in the reasoning logic of these NL2SQL models limits their application in high-risk scenarios or complex database requirements (spider2, ). To address these limitations, we explore the application of reinforcement learning to enhance the reasoning and generalization capabilities of NL2SQL models. By dynamically adjusting the model’s decision-making strategy through interaction with the environment, we aim to improve its performance in complex database scenarios and ensure better domain adaptation and interpretability.
Reinforcement Learning for LLM Reasoning.
Recent research on reinforcement learning for LLMs has increasingly concentrated on enhancing reasoning capabilities and optimizing interaction with the external environment to improve proficiency in complex, multi-step reasoning tasks (wei2022chain, ; plaat2024reasoning, ). In recent months, DeepSeek-R1 (deepseekr1, ) have illustrated the effectiveness of group relative policy optimization (GRPO), a novel and efficient reinforcement algorithm in cultivating robust reasoning behaviors in LLMs. By focusing exclusively on the correctness of the final answer, GRPO facilitates the internalization of intermediate reasoning steps without the necessity of explicit supervision, achieving notable success across various domains, including mathematics (shao2024deepseekmath, ), finance (finr1, ), programming (pan2025metaspatial, ), user interface (uir1, ; guir1, ) and vision tasks (visionr1, ). Concurrently, some studies have explored integrating external tools, such as retrieval (deepretrieval, ) and search engines (searchr1, ), into reinforcement learning frameworks for LLM reasoning, primarily focusing on training models to effectively employ a single, general-purpose tool to support specific reasoning chains. In light of these advancements, we aim to extend LLM reasoning capabilities to the NL2SQL task by adopting a reinforcement learning approach that dynamically adjusts the model’s decision-making strategy through interaction with the environment, thereby enhancing its ability to manage complex semantics and improve generalization and domain adaptation capabilities.
@@ 5 Limitations
Supported Database Dialect.
The current iteration of SQL-R1 has primarily been trained and evaluated on datasets that utilize the SQLite dialect. However, real-world databases often encompass various database dialects (e.g., Snowflake and DuckDB). Investigating and enhancing the generalization capabilities of NL2SQL models across these diverse database dialects presents a significant and valuable challenge for future research directions.
Experiment.
We conduct experiments on Qwen2.5-Coder series models. However, due to the rapid advancements in this field, our evaluation is limited to the capabilities of these representative LLMs and we can not cover a broader range of new LLMs available currently (e.g., Llama4). We provide an initial test result of more models on Section B.2, and we will keep extending our work on many more base LLMs.
@@ Appendix A Detailed Case Study
Appendix B Additional Results and Analysis
B.1 Performance on Other NL2SQL Benchmarks
Table 4 demonstrates the comparison of execution accuracy on Spider-DK, Spider-Syn and Spider-Realistic dataset. Notably, SQL-R1 exhibits superior performance relative to above NL2SQL solutions, which further prove the generalization and robustness of a under different database and query requirements.
Table 4: Execution accuracy (%) on Spider-DK, Spider-Syn, Spider-Realistic benchmark.
NL2SQL Method Base Model Spider-DK Spider-Syn Spider-Realistic
SENSE [55] CodeLlama-7B 77.9 72.6 82.7
ROUTE [56] Llama3-8B 74.6 77.4 80.9
SQL-o1 [30] Llama3-8B 78.7 72.6 82.7
OmniSQL [22] Qwen2.5-Coder-7B 77.8 69.6 78.0
SQL-PaLM [57] PaLM-2 67.5 70.9 77.4
PURPLE [58] GPT-4 75.3 74.0 79.9
SQL-R1 (Ours) Qwen2.5-Coder-3B 70.5 66.4 71.5
SQL-R1 (Ours) Qwen2.5-Coder-7B 78.1 76.7 83.3
SQL-R1 (Ours) Qwen2.5-Coder-14B 79.3 78.5 86.2
B.2 Analysis and Comparison of Base LLMs
As shown in Table 5, we found that smaller models benefit significantly more from RL training than their larger counterparts. Specifically, SQL-R1 trained on Qwen2.5-Coder-3B and Qwen2.5-Coder-7B demonstrated substantial accuracy improvements on both the Spider and BIRD benchmarks. In contrast, the Qwen2.5-Coder-14B model exhibited a relatively minor improvement. These results suggest that RL is particularly effective for smaller models, likely due to their greater capacity for performance enhancement through targeted training. This observation leads to two critical insights: first, the potential for deploying RL on smaller base models could yield significant performance gains with higher cost efficiency, enhancing their practicality for real-world applications; second, larger-scale models, such as the 14B variant, may necessitate more extensive training data and advanced training strategies to realize meaningful improvements. Future research could thus focus on optimizing RL methodologies for smaller models to maximize their potential, while also exploring sophisticated training techniques and larger datasets to leverage the capabilities of more complex LLMs fully.
Table 5: Execution accuracy (%) of different base LLMs on Spider and BIRD benchmark.
Base Model Spider (Dev) Spider (Test) Bird (Dev)
Qwen2.5-Coder-3B 77.0 77.2 50.5
Qwen2.5-Coder-7B 77.4 79.4 58.2
Qwen2.5-Coder-14B 87.0 88.0 66.1
Qwen3-8B (Thinking Mode) - - 50.8
Qwen3-14B (Thinking Mode) - - 51.8
SQL-R1  + Qwen2.5-Coder-3B 78.1 78.9 54.6
SQL-R1  + Qwen2.5-Coder-7B 87.6 88.7 63.1
SQL-R1  + Qwen2.5-Coder-14B 86.7 88.1 67.1
B.3 Ablation Study of SQL Candidate Rollouts in Reasoning
To assess the efficacy of the self-consistency in the inference, we conducted ablation studies on the Qwen2.5-Coder-3B and Qwen2.5-Coder-7B base models, aiming to determine the optimal number of SQL candidate rollouts generated under varied temperature settings designed to enhance diversity. The results illustrated in Figure 9 reveal that setting the number of SQL candidates to 8 strikes the ideal balance between diversity and computational efficiency. In the case of the 3B model, configurations with fewer than eight candidates resulted in diminished execution accuracy due to insufficient diversity, whereas exceeding 8 candidates led to a plateau in accuracy alongside increased computational overhead. The 7B model demonstrated relatively stable accuracy across different candidate counts but achieved its highest accuracy at the same optimal setting. These findings substantiate that the parameter configuration of 8 SQL candidates and a temperature setting of 0.8 effectively enhances SQL-R1’s performance, successfully navigating the trade-off between candidate diversity and computational cost, thereby reinforcing the model’s robustness in SQL query generatio
Figure 9: Performance of different scale of SQL candidate rollouts on BIRD-Dev dataset.
B.4 Ablation Study of Database Value Retrieval for RL Training
We conducted an ablation study to evaluate the effects of integrating database values during the RL training. As illustrated in the Table 6 , incorporating database values did not yield a statistically significant improvement in the model’s overall performance. However, this integration resulted in an increased input sequence length, which consequently diminished training efficiency. These findings substantiate our decision to forgo the inclusion of representative value annotations during the training phase. This strategic choice enables the model to concentrate more effectively on schema linking, thereby enhancing its exploratory capabilities during the RL phase and ultimately leading to improved reasoning and generalization skills.
Table 6: Ablation study of database value retrieval for RL training on BIRD-Dev dataset.
Base Model Use Database Values Accuracy (%)
Qwen2.5-Coder-7B - 58.2
SQL-R1  + Qwen2.5-Coder-7B
61.9
SQL-R1  + Qwen2.5-Coder-7B 63.1
B.5 Visualization of Training Process
Figure 10 illustrates the progression of reward and response length throughout the training process on Qwen2.5-Coder-3B, Qwen2.5-Coder-7B and Qwen2.5-Coder-14B models.
(a) Qwen2.5-Coder-3B
(b) Qwen2.5-Coder-7B
(c) Qwen2.5-Coder-14B
Figure 10: Visualization of training process on different scale base models.
Appendix C Prompt Templates